OpenStudy (anonymous):

could use some help to figure out the answer.it says use the net to find the surface area of the cylinder,give your answer in terms of pi. it has a square with 2 circles one on to and one on the bottom,the bottom circle has 4yd on the side of the square it has 14yd.

OpenStudy (anonymous):

 surface area of a circle = pi*r^2 if you take the top and bottom off of a cylinder and cut it then it will form a rectangle, the length of the rectangle is given, the width is the circumference of the circle which is: 2piR or dR add the area of the unrolled cylinder (aka rectangle) and the 2 circles to get total area.

OpenStudy (anonymous):

still some what confused not a big fan of Geometry.. could you show me how to break it down with the circle 4yd and the side cylinder 12yd..just so i have a better understanding..thank you

OpenStudy (anonymous):



OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the second picture?

OpenStudy (anonymous):

yes 4yd and the first on 12yd how do i solve it in terms of pi..sorry i am not very good at this

OpenStudy (anonymous):

ok so 4 yds from the middle to the edge? and 12 yards tall?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(your original instruction said 14 yards, I will use 12

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok for the area of a circle we use pi * r^2 pi = 3.14, 4yds squared is 16 yards, =16pi squared or 50.24yds squared. you have two of those circles for the cylinder that is now a rectangle: the height is 12yds, the width is equal to the circumference of the circle. circumference is diameter * pi 8*pi (times the height of 12yd) = an area of 96pi so total surface area: 96pi+16pi+16pi = 128pi squared or 402.1238597 yds squared

OpenStudy (anonymous):

wow that is amazing..sorry i have never did Geometry before..thank you so much for helping me figure this out..when they pi the only pi i know is the one you eat =) if i can get past these last 2 post i will be done with school October 31st..thank you so much.

OpenStudy (anonymous):

you bet, good luck.